# Data-driven geometric parameter optimization for PD-GMRES

### v1

@@ 1 Introduction
The structure of this work is as follows. In Section 2 the PD-GMRES algorithm from [4] is reviewed. We further show some theoretical results concerning the convergence behavior depending on specific parameter ranges. In Section 3 we introduce the quadtree-based geometric approach and adapt it to the optimization of the PD-GMRES parameters. This includes the use of averaged heuristic runtime estimates and the grouping of PD-GMRES parameters in pairs to apply our optimization approach, reducing a five-dimensional optimization problem to two dimensions. Section 4 comprises numerical experiments in which we investigate the application of our geometric approach to optimize PD-GMRES parameters for a set of training matrices. The optimization procedure yields an optimized PD-GMRES that exhibits very good convergence rates for a significant number of matrices. In Section 4.6 we propose an extension of the PD-GMRES algorithm using an additional parameter $m_{max}$ and showcase that this extended algorithm can provide better performance for a matrix for which the rate of convergence is very slow. We end with the concluding Section 5.
@@ 2 The PD-GMRES algorithm
[…]
@@
As the control equation (2.1) can cause the restart parameter $m_{j+1}$ to become too small to be useful or even negative, an additional parameter $m_{\min}\in{\mathbb{N}}$ is introduced as the minimal allowed restart parameter. If $m_{j+1}$ falls below $m_{\min}$ it resets to $m_{\text{init}}+c\cdot m_{\text{step}}$ (as simply resetting to $m_{\text{init}}$ would create a risk of stagnation). Here, $m_{\text{step}}\in{\mathbb{N}}$ is an additional parameter which determines by how much this reset value increases and $c$ is a counter for the resets initialized at 00.
@@
For $\alpha_{p}>0,\alpha_{d}>0$, the restart parameter increases directly in the case of stagnation, similarly guaranteeing convergence as the restart parameter will reach $n$. In this parameter range it is possible for the proportional and derivative terms in the control equation (2.1) to sum up to 00, resulting in $m_{j+1}=m_{j}$. In that case, the ratio $\frac{\|r_{j-2}\|}{\|r_{j}\|}$ can be directly determined from $\alpha_{p}$ and $\alpha_{d}$.
@@
It must be noted that the case that the restart parameter does not change is not directly addressed by Proposition 2.3, as it is possible for $m_{j+1}=m_{j}$ while $10<P_{j}(\alpha_{p})+D_{j}(\alpha_{d})<$ 1. In that case the floor function in the control equation (2.1) rounds down, and the rate of convergence is slightly worse than described in Proposition 2.3. Because of this, the systematic error caused by the rounding in the controller becomes undesirable. To reduce this error, we adjust the control algorithm by storing the remainder of the rounding operation and re-adding it during the next iteration. As the rounding was actually used in Proposition 2.1, the modification of storing and re-adding the rounding error is only applied if $\alpha_{p}>0$.
@@ 3 Geometric optimization approach
For the purpose of determining the PD-GMRES parameters $m_{\text{init}}$, $m_{\min}$, $m_{\text{step}}$, $\alpha_{p}$, $\alpha_{d}$ that minimize the PD-GMRES runtime, we describe a general quadtree-based algorithm for finding minima in a bounded two-dimensional domain which we then adapt to optimize the PD-GMRES parameters in pairs.
@@ 3.1 The quadtree method
0:  $f:D\rightarrow{\mathbb{R}},\ d\in{\mathbb{N}},\ \textbf{criterion C},\ \phi:Q\rightarrow D$
0:  $Q\in{\mathbb{R}}^{2^{d}\times 2^{d}}$
@@ 3.2 Application to PD-GMRES
In this section we apply the quadtree approach from Section 3.1 to minimize the runtime of PD-GMRES by optimizing the algorithm’s free parameters $m_{\text{init}}$, $m_{\min}$, $m_{\text{step}}$, $\alpha_{p}$, $\alpha_{d}$ in pairs.
@@ 3.2.1 A heuristic runtime estimate
As runtime measurements are volatile, in particular relative to the differences one can expect for slight parameter changes in PD-GMRES, we instead use a heuristic function which can deterministically calculate a runtime estimate. The function we use is
$\bar{h}(\tilde{r})=\sum_{\{j\in{\mathbb{N}}\ \!|\ \mathrm{Tol}_{\min}<\|\tilde{r}_{j}\|<\mathrm{Tol}_{\max}\}}h(m_{j})\frac{\log\left(\frac{\|\tilde{r}_{j-1}\|}{\|\tilde{r}_{j}\|}\right)}{\log\left(\frac{\|\tilde{r}_{k_{\min}}\|}{\|\tilde{r}_{k_{\max}}\|}\right)},$ (3.1)
where $\tilde{r}=(\tilde{r}_{j})$ is the sequence of the residual vectors of the iteration at each inner iteration of PD-GMRES (not just at each restart), and $\tilde{r}_{k_{\min}}$ and $\tilde{r}_{k_{\max}}$ are the residuals with the smallest and largest norm within $[\mathrm{Tol}_{\min},\mathrm{Tol}_{\max}]$ respectively. The function $h:{\mathbb{N}}\rightarrow{\mathbb{R}}_{+}$ gives a runtime estimate $h(m_{j})$ for one iteration of GMRES($m_{j}$), i.e., $m_{j}$ iterations of GMRES. We choose $h$ as a quadratic polynomial and determine the coefficients through numerical experiments by fitting runtime measurements for each matrix individually as each matrix exhibits different runtime costs. The choice of a quadratic polynomial is justified by the fact that the runtime complexity of performing $m$ steps of GMRES is $\mathcal{O}(m^{2})$ [16]. The function $\bar{h}$ (3.1) averages the individual heuristics $h$ over a range of residuals $\tilde{r}_{j}$. Specifically, we have
$\sum_{\{j\in{\mathbb{N}}\ \!|\ \mathrm{Tol}_{\min}<\|\tilde{r}_{j}\|<\mathrm{Tol}_{\max}\}}\frac{\log\left(\frac{\|\tilde{r}_{j-1}\|}{\|\tilde{r}_{j}\|}\right)}{\log\left(\frac{\|\tilde{r}_{k_{\min}}\|}{\|\tilde{r}_{k_{\max}}\|}\right)}=1,$
from which it follows that the heuristic runtime estimates $h(m_{j})$ for all residual norms between $\mathrm{Tol}_{\min}$ and $\mathrm{Tol}_{\max}$ are weighted in proportion to the improvement of the residual norm during the $j$-th step of PD-GMRES. This ensures that instead of optimizing for one specific tolerance we achieve good performance on average over a range of practical values.
@@
For the remainder of this paper when we refer to runtime, we use the averaged heuristic function $\bar{h}$ instead of actual runtime measurements.
@@ 4.1 Examples of PD-GMRES and quadtree behavior
Figure 4.1: Example quadtrees and PD-GMRES iteration behavior for the $(\alpha_{p}$, $\alpha_{d})$-space (top) and the $(m_{\min},m_{\text{step}})$-space (bottom).
@@ 4.2 Applying the optimization approach to training matrices
add20 $2395$ $13151$ $041.76e+$ 04 circuit simulation problem
cavity05 $1182$ $32632$ $059.18e+$ 05 computational fluid dynamics problem sequence
cavity10 $2597$ $76171$ $064.46e+$ 06 computational fluid dynamics problem sequence
cdde1 $961$ $4681$ $034.08e+$ 03 computational fluid dynamics problem sequence
circuit_2 $4510$ $21199$ $067.12e+$ 06 circuit simulation problem
ex40 $7740$ $456188$ $071.20e+$ 07 computational fluid dynamics problem
fpga_trans_01 $1220$ $7382$ $042.88e+$ 04 circuit simulation problem sequence
memplus $17758$ $99147$ $052.67e+$ 05 circuit simulation problem
orsirr_1 $1030$ $6858$ $051.67e+$ 05 computational fluid dynamics problem
orsreg_1 $2205$ $14133$ $041.54e+$ 04 computational fluid dynamics problem
pde2961 $2961$ $14585$ $029.49e+$ 02 2D/3D problem
raefsky1 $3242$ $293409$ $043.16e+$ 04 computational fluid dynamics problem sequence
raefsky2 $3242$ $293551$ $041.08e+$ 04 subsequent computational fluid dynamics problem
rdb2048 $2048$ $12032$ $031.96e+$ 03 computational fluid dynamics problem
sherman1 $1000$ $3750$ $042.26e+$ 04 computational fluid dynamics problem
sherman3 $5005$ $20033$ $166.90e+$ 16 computational fluid dynamics problem
sherman4 $1104$ $3786$ $037.16e+$ 03 computational fluid dynamics problem
sherman5 $3312$ $20793$ $053.90e+$ 05 computational fluid dynamics problem
steam2 $600$ $5660$ $063.55e+$ 06 computational fluid dynamics problem
wang2 $2903$ $19093$ $043.18e+$ 04 subsequent semiconductor device problem
wang4 $26068$ $177196$ $044.91e+$ 04 semiconductor device problem
watt_1 $1856$ $11360$ $095.38e+$ 09 computational fluid dynamics problem
young3c $841$ $3988$ $041.15e+$ 04 acoustics problem
@@
Final parameters for the optimized PD-GMRES are $m_{\text{init}}=10,m_{\min}=3,m_{\text{step}}=10,\alpha_{p}=-0.625,\alpha_{d}=4.375$. The top row depicts the $(\alpha_{p},\alpha_{d})$-space, and the bottom row the $(m_{\min},m_{\text{step}})$-space.
@@
$m_{\text{init}}=10,m_{\min}=3,m_{\text{step}}=10,\alpha_{p}=-0.625,\alpha_{d}=4.375$, geometric mean = $1$
@@ 4.3 Evaluating performance of optimized PD-GMRES on test matrices
ACTIVSg2000 $4000$ $28505$ $071.38e+$ 07 power network problem
CAG_mat1916 $1916$ $195985$ $081.37e+$ 08 combinatorial problem
Chem97ZtZ $2541$ $7361$ $024.63e+$ 02 statistical/mathematical problem
add32 $4960$ $19848$ $022.14e+$ 02 circuit simulation problem
adder_dcop_01 $1813$ $11156$ $081.30e+$ 08 subsequent circuit simulation problem
bcsstk34 $588$ $21418$ $045.09e+$ 04 structural problem
bfwa398 $398$ $3678$ $037.58e+$ 03 electromagnetics problem
cavity01 $317$ $7280$ $046.91e+$ 04 computational fluid dynamics problem sequence
cdde2 $961$ $4681$ $021.29e+$ 02 subsequent computational fluid dynamics problem
circuit_1 $2624$ $35823$ $053.27e+$ 05 circuit simulation problem
ex37 $3565$ $67591$ $022.26e+$ 02 computational fluid dynamics problem
fpga_trans_02 $1220$ $7382$ $042.91e+$ 04 subsequent circuit simulation problem
linverse $11999$ $95977$ $038.24e+$ 03 statistical/mathematical problem
msc01440 $1440$ $44998$ $067.00e+$ 06 structural problem
nasa1824 $1824$ $39208$ $066.63e+$ 06 structural problem
nasa2146 $2146$ $72250$ $033.97e+$ 03 structural problem
nasa2910 $2910$ $174296$ $071.76e+$ 07 structural problem
orsirr_2 $886$ $5970$ $051.67e+$ 05 computational fluid dynamics problem
pde900 $900$ $4380$ $022.93e+$ 02 2D/3D problem
pores_3 $532$ $3474$ $056.64e+$ 05 computational fluid dynamics problem
rdb5000 $5000$ $29600$ $034.30e+$ 03 computational fluid dynamics problem
steam1 $240$ $2248$ $072.99e+$ 07 computational fluid dynamics problem
wang1 $2903$ $19093$ $041.69e+$ 04 semiconductor device problem sequence
wang3 $26064$ $177168$ $041.07e+$ 04 semiconductor device problem
watt_2 $1856$ $11550$ $121.37e+$ 12 computational fluid dynamics problem
young1c $841$ $4089$ $029.85e+$ 02 acoustics problem

### v2

@@ 1 Introduction
In summary, the main contributions of this paper are the following.
  • 1.
    We present a general quadtree-based method for evaluation and optimization of functions on rectangular two-dimensional domains.
  • 2.
    We employ a heuristic runtime estimate for GMRES-based iterative solvers to avoid volatile runtime measurements.
  • 3.
    We develop an optimization procedure based on the quadtree method in which PD-GMRES parameters are grouped and optimized alternately to improve PD-GMRES performance. This method can be applied to arbitrary matrix sets to find optimized parameters.
  • 4.
    We apply our optimization procedure to a general set of training matrices and validate the parameters with a set of test matrices. From this we make recommendations for generic PD-GMRES parameter sets which can be used as default values when utilizing PD-GMRES.
The structure of this work is as follows. In Section 2 the PD-GMRES algorithm from [4] is reviewed. We further show some theoretical results concerning the convergence behavior depending on specific parameter ranges. In Section 3 we introduce the quadtree-based geometric approach and adapt it to the optimization of the PD-GMRES parameters. This includes the use of averaged heuristic runtime estimates and the grouping of PD-GMRES parameters in pairs to apply our optimization approach, reducing a five-dimensional optimization problem to two dimensions. Section 4 comprises numerical experiments in which we investigate the application of our geometric approach to optimize PD-GMRES parameters for a set of training matrices. The optimization procedure yields an optimized PD-GMRES that exhibits very good convergence rates for a significant number of matrices. In Section 4.7 we propose an extension of the PD-GMRES algorithm using an additional parameter $m_{\max}$ and showcase that this extended algorithm can provide better performance for a matrix for which the rate of convergence is very slow. We end with the concluding Section 5.
@@ 2 The PD-GMRES algorithm
PD-GMRES has been introduced in [4] as an adaptive extension of GMRES($m$). Starting with an initial vector $x_{0}$ to (1.1) and corresponding initial residual $r_{0}=Mx_{0}-b$ as well as an initial restart parameter $m_{1}=m_{\text{init}}\in{\mathbb{N}}$, PD-GMRES calls single iterations of GMRES($m_{j}$) which provide $x_{j}$ and the corresponding residual $r_{j}$ for the next restart. The next restart parameter $m_{j+1}$ is determined by the Proportional-Derivative (PD) controller based on the residual norms of the iteration after the previous two restarts (Algorithm 2.1), and GMRES($m_{j+1}$) is then executed on the previous iteration vector $x_{j}$ to get the next iteration vector $x_{j+1}$. The main PD-control equation is given by
@@
As the control equation (2.1) can cause the restart parameter $m_{j+1}$ to become too small to be useful or even negative, an additional parameter $m_{\min}\in{\mathbb{N}}$ is introduced as the minimal allowed restart parameter. If $m_{j+1}$ falls below $m_{\min}$ it resets to $m_{\text{init}}+c\cdot m_{\text{step}}$ (as simply resetting to $m_{\text{init}}$ would create a risk of stagnation). Here, $m_{\text{step}}\in{\mathbb{N}}$ is an additional parameter which determines by how much this reset value increases and $c$ is a counter for the resets initialized at 0.
@@
For $\alpha_{p}>0,\alpha_{d}>0$, the restart parameter increases directly in the case of stagnation, similarly guaranteeing convergence as the restart parameter will reach $n$. In this parameter range it is possible for the proportional and derivative terms in the control equation (2.1) to sum up to 0, resulting in $m_{j+1}=m_{j}$. In that case, the ratio $\frac{\|r_{j-2}\|}{\|r_{j}\|}$ can be directly determined from $\alpha_{p}$ and $\alpha_{d}$.
@@
It must be noted that the case that the restart parameter does not change is not directly addressed by Proposition 2.3, as it is possible for $m_{j+1}=m_{j}$ while $0<P_{j}(\alpha_{p})+D_{j}(\alpha_{d})<1$. In that case the floor function in the control equation (2.1) rounds down, and the rate of convergence is slightly worse than described in Proposition 2.3. Because of this, the systematic error caused by the rounding in the controller becomes undesirable. To reduce this error, we adjust the control algorithm by storing the remainder of the rounding operation and re-adding it during the next iteration. As the rounding was actually used in Proposition 2.1, the modification of storing and re-adding the rounding error is only applied if $\alpha_{p}>0$.
@@ 3 Geometric optimization approach
For the purpose of determining the PD-GMRES parameters $_{\text{init}}, mm$$_{\min}$, $m_{\text{step}}$, $\alpha_{p}$, $\alpha_{d}$ that minimize the PD-GMRES runtime, we describe a general quadtree-based algorithm for finding minima in a bounded two-dimensional domain which we then adapt to optimize the PD-GMRES parameters in pairs.
@@ 3.1 The quadtree method
[…]
@@ 3.2 Application to PD-GMRES
In this section we apply the quadtree approach from Section 3.1 to minimize the runtime of PD-GMRES by optimizing the algorithm’s free parameters $_{\text{init}}, mm$$_{\min}$, $m_{\text{step}}$, $\alpha_{p}$, $\alpha_{d}$ in pairs.
@@ 3.2.1 A heuristic runtime estimate
As runtime measurements are volatile, in particular relative to the differences one can expect for slight parameter changes in PD-GMRES, we instead use a heuristic function which can deterministically calculate a runtime estimate. A PD-GMRES iteration consists of individual iterations of GMRES($m_{k}$) with restart parameters $(m_{k})$. As a basis for our estimate we use $\sum_{k}h(m_{k})$, where the function $h:{\mathbb{N}}\rightarrow{\mathbb{R}}_{+}$ gives a runtime estimate $h(m_{k})$ for one iteration of GMRES($m_{k}$). We choose $h$ as a quadratic polynomial which is justified by the fact that the runtime complexity of performing $m$ steps of GMRES is $\mathcal{O}(m^{2})$ [16]. To determine $h$ we perform runtime measurements of single iterations of GMRES($m$) for multiple restart parameters $m$. We then calculate the coefficients of $h$ by least-squares fitting a quadratic polynomial to the measurement data. As our heuristic runtime estimate for a PD-GMRES iteration we use
$\bar{h}(\tilde{r})=\sum_{\{j\in{\mathbb{N}}\ \!|\ \mathrm{Tol}_{\max}>\|\tilde{r}_{j}\|>\mathrm{Tol}_{\min}\}}\left(\sum_{k}h\big((\tilde{m}_{j})_{k}\big)\right)\frac{\log\left(\frac{\|\tilde{r}_{j-1}\|}{\|\tilde{r}_{j}\|}\right)}{\log\left(\frac{\|\tilde{r}_{j_{\max}}\|}{\|\tilde{r}_{j_{\min}}\|}\right)},$ (3.1)
where $\tilde{r}=(\tilde{r}_{j})$ is the sequence of the residual vectors at each inner iteration of PD-GMRES (not just at each restart), $\tilde{r}_{j_{\min}}$ and $\tilde{r}_{j_{\max}}$ are the residuals with the smallest and largest norm within $\mathrm{Tol}_{\min},\mathrm{Tol}_{\max}]$ respectively and $((\tilde{m}_{j})_{k})$ are the restart parameters for the specific PD-GMRES iteration up to the inner index $j$, yielding the residual $\tilde{r}_{j}$ (which means that $\sum_{k}((\tilde{m}_{j})_{k})=j$). The $((\tilde{m}_{j})_{k})$ are determined from the restart parameters $(m_{k})$ of the complete PD-GMRES iteration. The function $\bar{h}$ from (3.1) averages the individual heuristics $\sum_{k}h\big((\tilde{m}_{j})_{k}\big)$ for all residuals $\tilde{r}_{j}$ between $\tilde{r}_{j_{\max}}$ and $\tilde{r}_{j_{\min}}$. Specifically, we have
$\sum_{\{j\in{\mathbb{N}}\ \!|\ \mathrm{Tol}_{\max}>\|\tilde{r}_{j}\|>\mathrm{Tol}_{\min}\}}\frac{\log\left(\frac{\|\tilde{r}_{j-1}\|}{\|\tilde{r}_{j}\|}\right)}{\log\left(\frac{\|\tilde{r}_{j_{\max}}\|}{\|\tilde{r}_{j_{\min}}\|}\right)}=1,$
from which it follows that the heuristic runtime estimates for all residual norms between $\mathrm{Tol}_{\max}$ and $\mathrm{Tol}_{\min}$ are weighted in proportion to the improvement of the residual norm during the $j$-th step of PD-GMRES. This ensures that instead of optimizing for one specific tolerance we achieve good performance on average over a range of practical values.
@@
In subsequent analysis we consider multiple different matrices and therefore determine a different heuristic $h$ for each matrix individually, as the runtime costs naturally differ.
For the remainder of this paper when we refer to runtime, we use the averaged heuristic function $\bar{h}$ instead of actual runtime measurements, for PD-GMRES as well as GMRES($m$).
@@ 3.2.2 Parameter space considerations
Each repetition of steps 3 and 4 likely improves the set of optimized parameters. With each optimization cycle one can expect diminishing returns for the performance gain of the parameters, which makes this optimization procedure a practical approach towards optimized, albeit not optimal, parameters. The limited resolution of the quadtree construction prevents a guarantee that a global minimum in the parameter space is found and therefore the optimization procedure can not be expected to converge to a set of globally optimal parameters. However, as observed in the numerical experiments (cf. Figure 4.2), in practice it is sufficient to terminate the procedure after a few optimization cycles.
@@ 4.1 Examples of PD-GMRES and quadtree behavior
Figure 4.1: Example quadtrees and PD-GMRES iteration behavior for the $(\alpha_{p}$, $\alpha_{d})$-space (top) and the $(m_{\min},m_{\text{step}})$-space (bottom). From each quadtree we take two samples for the matrices steam2 and pde2961, taking the local minimum for each (cyan for steam2 and green for pde2961) and the average minimum found accross all matrices (magenta for steam2 and red for pde2961). For these samples the behavior of the restart parameter and residual norm during the respective PD-GMRES iteration are shown.
@@ 4.2 Applying the optimization approach to training matrices
add20 $2395$ $13151$ $1.76e+04$ circuit simulation problem
cavity05 $1182$ $32632$ $9.18e+05$ computational fluid dynamics problem sequence
cavity10 $2597$ $76171$ $4.46e+06$ computational fluid dynamics problem sequence
cdde1 $961$ $4681$ $4.08e+03$ computational fluid dynamics problem sequence
circuit_2 $4510$ $21199$ $7.12e+06$ circuit simulation problem
ex40 $7740$ $456188$ $1.20e+07$ computational fluid dynamics problem
fpga_trans_01 $1220$ $7382$ $2.88e+04$ circuit simulation problem sequence
memplus $17758$ $99147$ $2.67e+05$ circuit simulation problem
orsirr_1 $1030$ $6858$ $1.67e+05$ computational fluid dynamics problem
orsreg_1 $2205$ $14133$ $1.54e+04$ computational fluid dynamics problem
pde2961 $2961$ $14585$ $9.49e+02$ 2D/3D problem
raefsky1 $3242$ $293409$ $3.16e+04$ computational fluid dynamics problem sequence
raefsky2 $3242$ $293551$ $1.08e+04$ subsequent computational fluid dynamics problem
rdb2048 $2048$ $12032$ $1.96e+03$ computational fluid dynamics problem
sherman1 $1000$ $3750$ $2.26e+04$ computational fluid dynamics problem
sherman3 $5005$ $20033$ $6.90e+16$ computational fluid dynamics problem
sherman4 $1104$ $3786$ $7.16e+03$ computational fluid dynamics problem
sherman5 $3312$ $20793$ $3.90e+05$ computational fluid dynamics problem
steam2 $600$ $5660$ $3.55e+06$ computational fluid dynamics problem
wang2 $2903$ $19093$ $3.18e+04$ subsequent semiconductor device problem
wang4 $26068$ $177196$ $4.91e+04$ semiconductor device problem
watt_1 $1856$ $11360$ $5.38e+09$ computational fluid dynamics problem
young3c $841$ $3988$ $1.15e+04$ acoustics problem
@@
Final parameters for the optimized PD-GMRES are $m_{\text{init}}=10,m_{\min}=3,m_{\text{step}}=10,\alpha_{p}=-0.625,\alpha_{d}=4.375$. The top row depicts the $(\alpha_{p},\alpha_{d})$-space, and the bottom row the $(m_{\min},m_{\text{step}})$-space. The geometric means for the average matrix performance (cf. Section 3.2.3) after each optimization step are: 1, 0.9270, 0.9060, 0.8957, 0.8957, 0.89531,\ 0.9270,\ 0.9060,\ 0.8957,\ 0.8957,\ 0.8953. These values are normalized around the average performance after the first optimization step which already improves performance significantly over GMRES($m$). We observe that the gains in matrix performance decrease with further optimization steps and therefore we terminate the optimization procedure after three cycles.
@@
$m_{\text{init}}=10,m_{\min}=3,m_{\text{step}}=10,\alpha_{p}=-0.625,\alpha_{d}=4.375$, geometric mean = 11
@@ 4.3 Evaluating performance of optimized PD-GMRES on test matrices
ACTIVSg2000 $4000$ $28505$ $1.38e+07$ power network problem
CAG_mat1916 $1916$ $195985$ $1.37e+08$ combinatorial problem
Chem97ZtZ $2541$ $7361$ $4.63e+02$ statistical/mathematical problem
add32 $4960$ $19848$ $2.14e+02$ circuit simulation problem
adder_dcop_01 $1813$ $11156$ $1.30e+08$ subsequent circuit simulation problem
bcsstk34 $588$ $21418$ $5.09e+04$ structural problem
bfwa398 $398$ $3678$ $7.58e+03$ electromagnetics problem
cavity01 $317$ $7280$ $6.91e+04$ computational fluid dynamics problem sequence
cdde2 $961$ $4681$ $1.29e+02$ subsequent computational fluid dynamics problem
circuit_1 $2624$ $35823$ $3.27e+05$ circuit simulation problem
ex37 $3565$ $67591$ $2.26e+02$ computational fluid dynamics problem
fpga_trans_02 $1220$ $7382$ $2.91e+04$ subsequent circuit simulation problem
linverse $11999$ $95977$ $8.24e+03$ statistical/mathematical problem
msc01440 $1440$ $44998$ $7.00e+06$ structural problem
nasa1824 $1824$ $39208$ $6.63e+06$ structural problem
nasa2146 $2146$ $72250$ $3.97e+03$ structural problem
nasa2910 $2910$ $174296$ $1.76e+07$ structural problem
orsirr_2 $886$ $5970$ $1.67e+05$ computational fluid dynamics problem
pde900 $900$ $4380$ $2.93e+02$ 2D/3D problem
pores_3 $532$ $3474$ $6.64e+05$ computational fluid dynamics problem
rdb5000 $5000$ $29600$ $4.30e+03$ computational fluid dynamics problem
steam1 $240$ $2248$ $2.99e+07$ computational fluid dynamics problem
wang1 $2903$ $19093$ $1.69e+04$ semiconductor device problem sequence
wang3 $26064$ $177168$ $1.07e+04$ semiconductor device problem
watt_2 $1856$ $11550$ $1.37e+12$ computational fluid dynamics problem
young1c $841$ $4089$ $9.85e+02$ acoustics problem